if the teacher (mass 85kg) was walking instead of running and had a momentum of 350kgm/s while doing so,what was their walking velocity?
1
Expert's answer
2012-03-13T07:25:55-0400
Let's V be the teacher's walking velocity. We know that P = M*V, so

V = P/M = (350kg*m/s) / (85kg) = 70/17 m/s ≈ 4.1176 m/s.